# Eating Speed Measurement Using Wrist-Worn IMU Sensors in Free-Living Environments

### v1

@@ 3.2.2 FD-II Dataset
The FD-II dataset contains 27 days of IMU data from 27 participants. The experiment protocol was the same as the FD-I dataset. However, in this dataset, only meal sessions were recorded by cameras (Some videos were collected by participants’ own smartphones). All other eating/drinking gestures outside of meal sessions were not recorded. Therefore, the ground truth information only contains the bite information during meals and the meal boundaries. The FD-II dataset is considered as a hold-out dataset, which contains 48 meals with 2,723 eating gestures (including four eating styles) over a total duration of 262 h.
3.3 Datasets for Training-Only
The FD-I and FD-II datasets are highly unbalanced, with the target classes (eating and drinking) being the minority. To further include more target data, we included two datasets containing IMU data collected in meal sessions, specially, the self-collected meal-only (MO) dataset and the external OREBA dataset [33] as part of the training set.
@@ 3.3.1 Meal-Only Dataset
The MO dataset contains 48 meal sessions from 48 participants, with a total of 2,894 eating and 763 drinking activities. It should be noted that part of this dataset was collected together with our previous Eat-Radar project [12] and there is no participant overlap between MO and FD datasets.
@@ 3.3.2 Public OREBA Dataset
The OREBA dataset [33] contains 100 meal sessions data from 100 participants, with 4,496 eating and 406 drinking gestures. The data were collected from both hands using two IMU wristbands. It should be noted that the coordinate system (direction of $x$, $y$, and $z$ axis) of the sensor used in OREBA is the same as ours, allowing us to integrate this data into the training set.
TABLE I: Full Day Datasets Statistics
@@ 3.4.1 Bite Annotation
Video recordings were viewed to annotate bites via ELAN [34]. The data were labeled into 3 classes, eating gesture, drinking gesture, and others. Three trained annotators labeled the datasets, with each annotator assigned to a specific portion of the dataset. The first author rechecked the annotation and made corrections when necessary. All annotators followed the same annotation instructions.
@@ 3.4.2 Eating Episodes Annotation
After annotating all the eating/drinking gesture from all-day data, the first eating gesture in an eating episode signifies the beginning boundary of the eating episode, and the last eating gesture in an eating episode is considered the ending boundary of the episode.
In free-living environments, it is normal for people to eat snacks outside meal sessions. However, snack eating exhibits high variability compared to meal eating. Some snack eating has frequent bites in very short duration, while other snack eating occurs over a longer period with a very low frequency (only one bite in several minutes). This variability makes the detection of snack session difficult; therefore, we focus on eating episodes that last at least 3 min. Snacking sessions with duration less than 3 min were neglected in this step.
@@ 3.4.3 Ground Truth Eating Speed
After annotating both eating gestures and eating episodes, we derived the number of bites in each eating episode and the duration of each respective eating episode. Consequently, the ground truth eating speed is obtained by computing the ratio of the number of bites in the eating episode and the time taken to finish the episode. The unit of eating speed is bites per minute (bite/min).
Figure 2: The quantity ratio of eating and drinking gestures using only dominant hand for each participant on FD-I dataset. A ratio with 0.5 means half number of eating gestures are from the dominant hand, the others are from the non-dominant hand, a ratio with -1 means all drinking gestures are from the dominant hand.
@@ 3.5 Data Preprocessing
Existing approaches in free-living environments typically process IMU data from the dominant hand, assuming that people only use one hand for daily food intake. However, in real-life, both hands can be used for eating food or drinking water, as illustrated in Fig. 2. Considering the IMU waveform for eating with left and right hand differs, we applied the two-hand combination method that combines hand mirroring and temporal concatenation to process two-hand IMU data, which was validated in our previous study [35]. The hand mirroring method has been applied in multiple IMU-based bite detection studies when the participant is left hand dominant, which involves flipping the direction of $a_{x}$ (in accelerometer), $g_{y}$ and $g_{z}$ (in gyroscope). The hand-mirrored left hand IMU data were than concatenated after right hand data. Hence, the preprocessed data have 6 channels, and the data length of each recording is doubled, as shown in Fig. 3. Meanwhile, to reduce the computation cost, the data were downsampled to 16 Hz.
@@ 3.6.2 MHA Module
The MHA module is an essential component of the Transformer [29], showcasing considerable efficacy in capturing relationships within time sequence signals. Before being fed into the attention module, the positional encoding is added to the input feature map. The attention mechanism is considered as a mapping among the query ($Q=X_{1}W_{Q}$), key ($K=X_{1}W_{K}$), and value ($V=X_{1}W_{V}$), where $X_{1}\in\mathbb{R}^{T\times{d_{in}}}$ signifies the input of the module ($d_{in}=C_{m}$), $W_{Q}$,$W_{K}$ and $W_{V}\in\mathbb{R}^{d_{in}\times{d_{model}}}$, with $d_{in}$ representing the dimension of the output sequence from the TCN model, and $d_{model}$ denoting the dimension of the attention module. The formula of the attention calculation is presented as follows:
@@
where $W^{O}\in\mathbb{R}^{h\cdot d_{h}\times{d_{model}}}$, $W_{i}^{Q}$, $W_{i}^{K}$ and $W_{i}^{V}\in\mathbb{R}^{d_{model}\times{d_{h}}}$, and $d_{h}$ denotes the dimension of each individual head in the attention mechanism, here we keep $d_{h}=d_{model}/h$.
@@ 3.6.3 FCN Module
3.6.4 Loss Function
The model’s loss function comprises both classification and smoothing loss. Specifically, the classification loss is implemented through a cross-entropy loss, while the smoothing loss involves a truncated mean squared error (MSE) calculated over sample-wise log-probabilities. Details on the integrated loss function can be referenced in [27].
@@ 3.6.5 Implementation Details
The TCN-MHA model is implemented within the PyTorch framework. Following each dilated layer, a dropout rate of 30% is applied. Based on experiments, the TCN module is constructed with a total of 9 layers, with each layer comprising 64 kernels. For the MHA module, 8 attention heads are employed, with each head characterized by a dimension of 16, resulting in a model dimension of 128 ($d_{model}$=128). The first layer of the FCN has 64 neurons with ReLU activation, the last layer contains 3 neurons with Softmax activation. For model training, an Adam optimizer is utilized with a learning rate set to 0.0005. It is important to note that the model exhibits a temporal lag in its predictions due to its non-causal nature. This temporal delay can be quantified by dividing half of the receptive field by the sampling frequency (0.5$\times$1023/16=32s). The window length of input data is set to 60 s accordingly.
All experiments for training, validation, and testing were carried out on a computational node equipped with an Intel 18-core Xeon Gold 6140 CPUs@2.3 GHz (Skylake) with 5 GB RAM per core, and two NVIDIA P100-SXM2-16 GB GPUs provided by Vlaams Supercomputer Centrum (VSC)  [^4] See https://www.vscentrum.be/.
@@ 3.7 Bite Detection
Figure 5: The DBSCAN clustering example for eating episode detection.
@@ 3.8 Eating Episode Detection
In bite detection step, to achieve data uniformity, the data from the left hand is hand mirrored and then temporally concatenated after right hand data. Prior to meal session detection, the output data is divided into to two sub-sequences (left hand and right hand), then an OR operation is applied to integrate bites from both hands. As eating episodes mainly involves eating gestures, all detected drinking gestures are removed in this step.
The predicted bite sequence is then clustered by 1D density-based spatial clustering of applications with noise algorithm (1D-DBSCAN) [25] to compose eating episodes. The DBSCAN identifies clusters based on the density of points. The function sklearn.cluster.DBSCAN is employed with an epsilon parameter set to 3 min and a minimum samples parameter set to 5. In our case, the distance between two bites is the temporal proximity between the bites. Subsequently, sparse bites are filtered as noise, while bites sequence with high density are clustered together to compose eating episodes, as shown in Fig. 5. After clustering, we follow the same operations in [6, 36] to merge very close eating episodes and remove very short episodes. Specifically, if the distance between two eating episodes is less than 3 min, they are merged into one episode. Additionally, if the duration of the episode is less than 3 min after merging, the eating episode is removed. Finally, we obtain a set of detected eating episode =$\mathcal{M}=\left\{m_{1},m_{2},...m_{Z}\right\}$, where each $m_{j}$ corresponds to the interval $[t_{j}^{l},t_{j}^{r}]$ ($1\leq{j}\leq{Z}$) that corresponds the left and right boundaries of $j$-th eating episode, $Z$ denotes the total number of detected eating episodes.
@@ 3.9 Eating Speed Estimation
After yielding eating episodes from the previous step, eating speeds are obtained by utilizing the previously detected bite set $\mathcal{B}$ and the eating episode set $\mathcal{M}$. The eating episode detection algorithm predicts the start point $t_{j}^{l}$ and end point $t_{j}^{r}$ of the $j$-th eating episode. If the detected bite $b_{i}$ falls within the interval of the eating episode, this bite is considered to belong to that eating episode. At the end, the number of bites divided by the duration of the eating episode gives the eating speed.
Figure 6: Examples of segment-wise evaluation. When the evaluated class is drinking gesture, eating gestures are categorized as other, and vice versa.
@@ 4.1.1 Evaluation on Bite Detection
The output of the proposed seq2seq model is point-wise multi-class prediction. As bite-related datasets are normally unbalanced, we choose to use the index Cohen Kappa [37] to represent the performance of point-wise classification. Although such results can indicate the performance of the model, it should be noted that the purpose of bite detection is to count the number of bites, whereas point-wise results are unable to reveal such information. To address this issue, we use a segment-wise evaluation method to evaluate the bite detection, which has been applied in previous study [12]. Fig. 6 shows examples of this evaluation. The evaluation method involves two steps. Firstly, the intersection over union (IoU) between each predicted bite and ground truth bite is calculated, as shown in Fig. 6 C1-C3. Secondly, the calculated IoU is compared to a selected threshold $k$ to determine segment-wise true positive (TP), false negative (FN) and false positive (FP). Subsequently, the segmental F1-score is calculated for each class (eating and drinking).
The segment-wise evaluation scheme allows for short temporal shifts between ground truth and prediction, which maybe caused by annotation variability. Meanwhile, it furnishes straightforward information including the number of detected bites. Furthermore, by adjusting the threshold $k$, we can evaluate not only the detection performance, but also the segmentation performance. In this study, two thresholds are selected as 0.1 and 0.5.
@@ 4.1.2 Evaluation on Eating Episode Detection
The evaluation of eating episode detection mainly focuses on two aspects, specifically, the detection performance (how many eating episodes are detected), and the segmentation performance (how well the boundaries of each eating episode are determined). Therefore, the aforementioned segment-wise evaluation method is also used for eating episode detection ($k=0.5$). Additionally, for each predicted eating episode, we utilize the IoU score to evaluate the segmentation performance.
@@ 4.2 Models for Benchmarking
To evaluate the efficacy of the proposed method, existing models for bite detection were chosen as comparative benchmarks. Specifically, the CNN-LSTM [6], ResNet-LSTM [21], and MS-TCN [32] were chosen. Additionally, the bidirectional type of LSTM layer (BiLSTM) was used to replace the LSTM layer to compose the CNN-BiLSTM and ResNet-BiLSTM models as extra models. All of these models can learn temporal context from time-series input data and have demonstrated their capability in human activity recognition. It should be noted that the output of these models are point-wise probabilities.
TABLE II: Bite Detection Performance on FD-I Dataset
@@
TABLE III: Eating Episode Detection and Eating Speed Performance on FD-I Dataset.
@@
Figure 7: The scatter plot for eating speed on FD-I dataset.
@@
TABLE IV: In-meal Bite Detection, Eating Episode Detection and Eating Speed Performance on FD-II Dataset
[…]
@@
FD-II CNN-LSTM 0.731 0.555 43 53 9 0.581 0.746 0.231 0.683
CNN-BiLSTM 0.783 0.622 47 35 5 0.701 0.786 0.173 0.860
ResNet-LSTM 0.752 0.584 48 42 4 0.676 0.809 0.201 0.780
ResNet-BiLSTM 0.798 0.650 49 26 4 0.766 0.827 0.142 0.910
(Hold-out) MS-TCN 0.814 0.636 47 22 5 0.777 0.831 0.155 0.886
TCN-MHA 0.820 0.651 48 7 4 0.897 0.841 0.146 0.924

### v2

@@ 3.2.2 FD-II Dataset
The FD-II dataset contains 27 days of IMU data from 27 participants. The experiment protocol was the same as the FD-I dataset. However, in this dataset, only meal sessions were recorded by cameras (Some videos were collected by participants’ own smartphones). All other eating/drinking gestures outside of meal sessions were not recorded. Therefore, the ground truth information only contains the bite information during meals and the meal boundaries. The FD-II dataset is considered as a holdout dataset, which contains 52 meals with 2,722 eating gestures (including four eating styles) over a total duration of 261.68 h.
3.3 In-Meal Datasets
The FD-I and FD-II datasets are highly unbalanced, with the target classes being the minority. To further include more target data, we included two datasets containing IMU data collected in meal sessions, specially, the self-collected meal-only (MO) dataset and the external OREBA dataset [35] as part of the training set. Additionally, the OREBA and Clemson [20] datasets were included for comparison analysis.
@@ 3.3.1 Meal-Only Dataset
The MO dataset contains 46 meal sessions from 46 participants, including 2,894 eating and 763 drinking gestures. It should be noted that part of this dataset was collected together with our Eat-Radar project [12] and there is no participant overlap between MO and FD datasets.
@@ 3.3.2 Public OREBA Dataset
The OREBA dataset [35] contains 100 meal sessions data from 100 participants, with 4,496 eating and 406 drinking gestures. The data were collected from both hands using two IMU wristbands. We downsampled the data from 64 Hz to 16 Hz. It should be noted that the coordinate system (direction of $x$, $y$, and $z$ axis) of the sensor used in OREBA is the same as ours, allowing us to integrate this data into the training set.
3.3.3 Public Clemson Dataset
The Clemson dataset [20] contains 488 meal sessions across 264 participants, with a total of 18,430 eating and 2,152 drinking gestures. The data were collected by placing a wristband IMU on the dominant hand and the sampling rate of data is 15 Hz.
Table 1: Full Day Datasets Statistics
@@ 3.3.3 Public Clemson Dataset
Figure 3: The pie chart for eating styles (a) and eating locations (b). The eating speeds represent the average speed of each category.
Figure 4: The violin plot of ground truth eating speed distribution among different datasets.
@@ 3.4.1 Bite Annotation
Video recordings were viewed to annotate bites via ELAN [36]. The data were labeled into 3 classes, eating gesture, drinking gesture, and others. Three trained annotators labeled the datasets, with each annotator assigned to a specific portion of the dataset. The first author rechecked the annotation and made corrections when necessary. All annotators followed the same annotation instructions. To synchronize the video and IMU data, the IMU data were divided into segments using the timestamps recorded at the start of each video, as described in Section III-B-1. For IMU segments corresponding with video recordings (state 1), the timelines were aligned by matching the IMU timestamps with those at the beginning of each video. The ELAN tool was used to import both the video and IMU data to verify the alignment and manually adjust it if there was a time shift (normally 1-2 s). Timing shift was identified by observing whether the IMU waveform response to the participants’ hand movements, transitioning from stillness to motion and vice versa, was delayed or advanced. For IMU segments without video recordings (state 2), labels of ’0’ were assigned directly. These labeled segments were then concatenated together, as shown in Fig. 2(b).
@@ 3.4.2 Eating Episodes Annotation
The first eating gesture in an eating episode signifies the beginning boundary of the eating episode, and the last eating gesture in an eating episode is considered the ending boundary of the episode. In free-living environments, it is normal for people to eat snacks outside meal sessions. However, snack eating exhibits high variability compared to meal eating. Some snack eating has frequent bites in very short duration, while other snack eating occurs over a longer period with a very low frequency (only one bite in several minutes). This variability makes the detection of snack session difficult; therefore, we focus on episodes that last at least 3 min. Snacking sessions with duration less than 3 min were neglected in this step. According to our definition, there are 457 episodes in Clemson and 101 episodes in OREBA.
@@ 3.4.3 Ground Truth Eating Speed
The ground truth eating speed is obtained by computing the ratio of the number of bites in the eating episode and the duration of the episode. The unit of eating speed is bites per minute (bite/min). Fig. 4 shows the distribution of eating speed in different datasets.
Figure 5: The quantity ratio of eating gestures using only dominant hand on FD-I dataset. A ratio with 0.5 means half number of eating gestures are from the dominant hand, the others are from the non-dominant hand.
@@ 3.5 Data Preprocessing
In daily life, both hands can be used for eating food, as illustrated in Fig. 5. Considering the IMU waveform for eating with left and right hand differs, we applied the two-hand combination method that combines hand mirroring and temporal concatenation to process two-hand IMU data, which was validated in our previous study [37]. The hand mirroring method has been applied in multiple IMU-based bite detection studies when the participant is left hand dominant, which involves flipping the direction of $a_{x}$ (in accelerometer), $g_{y}$ and $g_{z}$ (in gyroscope). The hand-mirrored left hand IMU data were than concatenated after right hand data. Hence, the preprocessed data have 6 channels, and the data length of each recording is doubled, as shown in Fig. 6. Meanwhile, to reduce the computation cost, the data were downsampled to 16 Hz.
@@ 3.6.2 MHA Module
The MHA module is an essential component of the Transformer [31], showcasing considerable efficacy in capturing relationships within time sequence signals. Before being fed into the attention module, the positional encoding is added to the input feature map. The attention mechanism is considered as a mapping among the query ($Q=X_{1}W_{Q}$), key ($K=X_{1}W_{K}$), and value ($V=X_{1}W_{V}$), where $X_{1}\in\mathbb{R}^{T\times{d_{in}}}$ signifies the input of the module ($d_{in}=C_{m}$), $W_{Q}$,$W_{K}$, and $W_{V}\in\mathbb{R}^{d_{in}\times{d_{model}}}$ are weight matrices used to transform the input data into Q, K, and V, respectively, with $d_{in}$ representing the dimension of the output from the TCN module, and $d_{model}$ denoting the dimension of the attention module. The formula of the attention calculation is presented as follows:
@@
where $W^{O}\in\mathbb{R}^{h\cdot d_{h}\times{d_{model}}}$, $W_{i}^{Q}$, $W_{i}^{K}$ and $W_{i}^{V}\in\mathbb{R}^{d_{model}\times{d_{h}}}$, and $d_{h}$ denotes the dimension of each individual head in the attention block, here $d_{h}=d_{model}/h$.
@@ 3.6.4 Implementation Details
The model’s loss function comprises both classification and smoothing loss. Specifically, the classification loss is a cross-entropy loss, while the smoothing loss involves a truncated mean squared error (MSE). Details on the integrated loss function can be referenced in [29]. Following each dilated layer in the TCN module, a dropout rate of 30% is applied. Based on experiments, the TCN module is constructed with a total of 9 layers, with each layer comprising 64 kernels. For the MHA module, 8 attention heads are employed, with each head characterized by a dimension of 16, resulting in a model dimension of 128 ($d_{model}$=128). The first layer of the FCN has 64 neurons with ReLU activation, the last layer contains 3 neurons with Softmax activation. For model training, an Adam optimizer is utilized with a learning rate set to 0.0005. It is important to note that the model exhibits a temporal lag in its predictions due to its non-causal nature. This temporal delay can be quantified by dividing half of the receptive field by the sampling frequency (0.5$\times$1023/16=32s). The window length of input data is set to 60 s accordingly. All experiments were carried out on two NVIDIA P100 GPUs provided by Vlaams Supercomputer Centrum (VSC)  [^4] See https://www.vscentrum.be/.
@@ 3.7 Bite Detection
Figure 8: The DBSCAN clustering example for eating episode detection. In the DBSCAN clustering step, arrows with different colors (yellow or green) represent different clusters.
@@ 3.8 Eating Episode Detection
In bite detection step, to achieve data uniformity, the data from the left hand is hand mirrored and then temporally concatenated after right hand data. Prior to meal session detection, the output data is divided into to two sub-sequences (left hand and right hand), then an OR operation is applied to integrate bites from both hands. As eating episodes mainly involves eating gestures, all detected drinking gestures are removed in this step. The predicted bite sequence is then clustered by 1D-DBSCAN [25] to compose eating episodes. The DBSCAN identifies clusters based on the density of points. The function sklearn.cluster.DBSCAN is employed with an epsilon parameter set to 3 min and a minimum samples parameter set to 5. In our case, the distance between two bites is the temporal proximity between the bites. Subsequently, sparse bites are filtered as noise, while bites sequence with high density are clustered together to compose eating episodes, as shown in Fig. 8. After clustering, we follow the same operations in [6, 38] to merge very close eating episodes and remove very short episodes. Specifically, if the distance between two eating episodes is less than 3 min, they are merged into one episode. Additionally, if the duration of the episode is less than 3 min after merging, the eating episode is removed. Finally, we obtain a set of detected eating episode =$\mathcal{M}=\left\{m_{1},m_{2},...m_{Z}\right\}$, where each $m_{j}$ corresponds to the interval $[t_{j}^{l},t_{j}^{r}]$ ($1\leq{j}\leq{Z}$) that corresponds the left and right boundaries of $j$-th eating episode, $Z$ denotes the total number of detected eating episodes.
@@ 3.9 Eating Speed Estimation
Eating speeds are obtained by utilizing the detected bite set $\mathcal{B}$ and eating episode set $\mathcal{M}$. The eating episode detection algorithm predicts the start point $t_{j}^{l}$ and end point $t_{j}^{r}$ of the $j$-th eating episode. If the detected bite $b_{i}$ falls within the interval of the eating episode, this bite is considered to belong to that eating episode. At the end, the number of bites divided by the duration of the eating episode gives the eating speed.
Figure 9: Segment-wise evaluation examples. If the target is drinking class, eating gestures are categorized as other, and vice versa.
@@ 4.1.1 Evaluation on Bite Detection
The output of the proposed seq2seq model is point-wise multi-class prediction. As bite-related datasets are normally unbalanced, we choose to use the index Cohen Kappa [39] to represent the performance of point-wise classification. Although such results can indicate the performance of the model, it should be noted that the purpose of bite detection is to count the number of bites, whereas point-wise results are unable to reveal such information. To address this issue, we use a segment-wise evaluation method to evaluate the bite detection, which has been applied in previous study [12]. Fig. 9 shows examples of this evaluation. The evaluation method involves two steps. Firstly, the intersection over union (IoU) between each predicted bite and ground truth bite is calculated, as shown in Fig. 9 C1-C3. Secondly, the calculated IoU is compared to a selected threshold $k$ to determine segment-wise true positive (TP), false negative (FN) and false positive (FP). Subsequently, the segmental F1-score is calculated for each class (eating and drinking). The segment-wise evaluation allows for short temporal shifts between ground truth and prediction, which maybe caused by annotation variability. Meanwhile, it furnishes straightforward information including the number of detected bites. Furthermore, by adjusting the threshold $k$, we evaluate not only the detection performance, but also segmentation performance. Two thresholds are selected as 0.1 and 0.5.
@@ 4.1.2 Evaluation on Eating Episode Detection
The evaluation of eating episode detection mainly focuses on two aspects, specifically, the detection performance, and the segmentation performance (how well the boundaries of each eating episode are determined). Therefore, the aforementioned segment-wise evaluation method is also used for eating episode detection ($k=0.5$). Additionally, for each predicted eating episode, we utilize the IoU score to evaluate the segmentation performance.
@@ 4.2 Models for Benchmarking
To evaluate the efficacy of the proposed method, existing models for bite detection were chosen as comparative benchmarks. Specifically, the CNN-LSTM [6], ResNet-LSTM [21], and MS-TCN [34] were chosen. Additionally, the bidirectional-LSTM layer (BiLSTM) was used to replace the LSTM layer to compose the CNN-BiLSTM and ResNet-BiLSTM models as extra models. It should be noted that the output of models are point-wise probabilities.
Table 2: Bite Detection Performance on FD-I Dataset
@@
Table 3: Eating Episode and Speed Detection on FD-I Dataset
@@
Figure 10: The scatter plot for eating speed on FD-I dataset. Black dots represent TP eating episodes, red dots are FNs, blue dots are FPs.
@@
Table 4: In-meal Bite Detection, Eating Episode Detection and Eating Speed Performance on FD-II Dataset
Data Model Eating Gesture F1-scorea Eating Episodes Detection Eating Speed # Paras (M) # FLOPs (G)
@@
FD-II CNN-LSTM 0.731 0.555 43 53 9 0.581 0.746 0.231 0.683 0.134 0.129
CNN-BiLSTM 0.783 0.622 47 35 5 0.701 0.786 0.173 0.860 0.241 0.233
ResNet-LSTM 0.752 0.584 48 42 4 0.676 0.809 0.201 0.780 3.078 2.955
ResNet-BiLSTM 0.798 0.650 49 26 3 0.772 0.827 0.142 0.910 3.415 3.280
(Holdout) MS-TCN 0.814 0.636 47 22 5 0.777 0.831 0.155 0.886 0.298 0.284
TCN-MHA 0.820 0.651 48 7 4 0.897 0.841 0.146 0.924 0.203 0.194
@@
Table 5: Results on OREBA and Clemson datasets using TCN-MHA
Valid Data Eating Gesture F1-score Eating Episodes Detection Eating Speed
$k=0.1$ $k=0.5$ TP FP FN F1 IoU MAPE PCC
7-fold OREBA 0.897 0.878 97 0 4 0.980 0.967 0.069 0.940
Clemson 0.894 0.835 445 2 12 0.985 0.931 0.097 0.913
hold OREBA 0.803 0.720 88 0 13 0.931 0.907 0.119 0.834
out Clemson 0.710 0.652 275 0 182 0.751 0.825 0.264 0.607
Table 6: Eating styles and locations performance using TCN-MHA on FD-I dataset
View Type Eating Episodes Detection Eating Speed
TP FN Recall IoU MAPE PCC
Cutlery Fork&knife 34 1 0.971 0.942 0.099 0.866
Chopsticks 6 1 0.857 0.961 0.136 0.982
Spoon 8 1 0.889 0.955 0.079 0.921
Hand 16 7 0.696 0.896 0.139 0.940
Location Home 12 4 0.750 0.922 0.100 0.956
Restaurant 31 0 1.000 0.943 0.113 0.923
Workspace 21 6 0.778 0.927 0.111 0.924